Jacob and Isaac pull firewood on a sled with a mass of 60kg to the right across a snow covered field. Jacob pulls with a force of 30N and Isaac pulls with a force of 20N and there is a force of friction of 15N.
1. What is the net force acting on the sled?
2. What is the acceleration of the sled?

Answer by fractalier(6550)

1) The net force is 30N + 20N - 15N = 35N to the right...friction almost always works in opposition to the motion...
2) Newton's Second Law gives a = F/m = 35N / 60kg = .583 m/s^2